### v1

# FlowKV: Enhancing Multi-Turn Conversational Coherence in LLMs via Isolated Key-Value Cache Management

Abstract
Large Language Models (LLMs) are increasingly deployed in multi-turn conversational applications, where the management of the Key-Value (KV) Cache presents a significant bottleneck. The linear growth of the KV Cache with dialogue history imposes substantial computational costs, and existing eviction strategies often degrade performance by repeatedly compressing early conversational context, leading to information loss and context forgetting. This paper introduces FlowKV, a novel multi-turn isolation mechanism for KV Cache management, which can be applied to any KV Cache compression method without training. FlowKV’s core innovation is a multi-turn isolation mechanism that preserves the accumulated compressed KV cache from past turns. Compression is then strategically applied only to the newly generated KV pairs of the latest completed turn, effectively preventing the re-compression of older context and thereby mitigating catastrophic forgetting. Our results demonstrate that FlowKV consistently and significantly outperforms baseline strategies in maintaining instruction-following accuracy and user preference retention from 10.90% to 75.40%, particularly in later conversational turns.
FlowKV: Enhancing Multi-Turn Conversational Coherence in LLMs via Isolated Key-Value Cache Management
Xiang LIU        Hong CHEN        Xuming HU        Xiaowen CHU The Hong Kong University of Science and Technology(Guangzhou) {xliu886,hchen763}@connect.hkust-gz.edu.cn {xuminghu, xwchu}@hkust-gz.edu.cn
**footnotetext: Equal Contribution.$\dagger$\$$\dagger$\$footnotetext: Corresponding Author.
1 Introduction
Large Language Models (LLMs) have achieved remarkable success in a wide range of natural language understanding and generation tasks, powering applications from open-domain dialogue systems to complex instruction following (Raffel et al., 2020; Brown et al., 2020; Chowdhery et al., 2022; Tay et al., 2022; Touvron et al., 2023a, b; Dubey et al., 2024; Guo et al., 2025). However, as these models are increasingly deployed in multi-turn conversational scenarios, new challenges emerge regarding their efficiency and ability to maintain coherent, contextually relevant responses over extended interactions.
Figure 1: Performance comparison of different KV Cache compression methods on multi-turn benchmark PrefEval. # SKV: SnapKV, # EA: ExpectedAttention.
The complexity of managing information flow in such multi-turn settings is evident in the attention patterns of LLMs. Figure 2 visualizes an attention heatmap from a representative 3-turn dialogue. The heatmap reveals several crucial characteristics: sustained attention to initial system prompts (SYS) throughout the conversation, strong local attention within each conversational turn, and, critically, significant cross-turn dependencies. For instance, responses not only attend to their immediate queries but also to previous queries and responses (e.g., the Turn 2 Response attending to T1Q and T1R). Furthermore, subsequent user queries (e.g., T2Q and T3Q) demonstrate an evolving understanding by referencing earlier queries and the system prompt. These intricate attention patterns, highlighting the model’s reliance on both recent and distant context, directly underscore the challenges in efficiently managing the conversational history.
Figure 2: Attention heatmap illustrating token-level focus during a 3-turn conversational interaction. The Y-axis represents query token positions, while the X-axis shows key token positions. Key observations (highlighted on the right) include: (1) Responses (e.g., Turn 1 Response) heavily focus on their corresponding query (e.g., T1Q) and the local context (local window). (2) As the dialogue progresses, responses (e.g., Turn 2 and 3 Responses) attend to an increasing span of historical context, including previous queries and responses (e.g., T1Q, T1R, T2Q, T2R). (3) Later queries (Turn 2 & 3 Queries) exhibit increased attention to both previous queries and the initial system prompt, indicating an evolving contextual understanding. These patterns underscore the complex, long-range dependencies managed by the attention mechanism in multi-turn dialogues.
This challenge of efficiently managing extensive conversational history, so critical for maintaining coherence as highlighted by the attention patterns, is primarily rooted in the operational characteristics of the KV Cache within the attention mechanism. While the KV Cache enables efficient autoregressive generation by storing intermediate representations, its memory footprint grows linearly with the length of the conversation history (Xiao et al., 2023). This not only imposes significant computational and storage costs but also limits the practical deployment of LLMs in real-world, resource-constrained environments.
To address these challenges, a variety of KV Cache eviction strategies have been proposed (Xiao et al., 2023; Zhang et al., 2023; Liu et al., 2025b; Li et al., 2024b; Liu et al., 2025a; Cai et al., 2024; Yang et al., 2024b). Despite their promise, the effectiveness of these methods in multi-turn dialogue settings remains underexplored. In particular, it is unclear how different compression strategies impact the model’s ability to follow instructions, retain user preferences, and maintain long-term dependencies across multiple conversational turns.
In this work, we conduct a comprehensive empirical study of KV Cache eviction strategies in multi-turn dialogue scenarios. We benchmark several state-of-the-art techniques—including SnapKV Li et al. (2024b), StreamingLLM Xiao et al. (2023), ExpectedAttention Jegou et al. (2024), and ChunkKV Liu et al. (2025b)—on two representative datasets: Multi-IF He et al. (2024), which evaluates instruction following across turns, and PrefEval Zhao et al. (2025), which assesses user preference retention. Figure 1 shows these methods’ performance on PrefEval, highlighting the KV cache compression method still have a large gap with the full KV cache. Furthermore, we introduce FlowKV, a multi-turn isolation mechanism designed to mitigate information loss and context forgetting under aggressive compression. This training-free approach is compatible with any existing KV Cache compression method and demonstrates significant improvements, for instance, boosting the average Instruction Following Rate (IFR) by over 20% on Multi-IF and increasing user preference adherence on PrefEval from as low as 10.90% to 75.40% with LLaMA-3.1-8B.
Our results demonstrate that FlowKV consistently outperforms baseline strategies, especially in later turns where context accumulation and compression effects are most pronounced. Through extensive quantitative and qualitative analyses, we provide new insights into the trade-offs between efficiency and conversational ability in LLMs, paving the way for more robust and scalable dialogue systems.
Figure 3: Illustration of KV cache dynamics across three management strategies in a two-turn conversational setting. Each turn consists of a system prompt (Sys Prompt), user query (Query), and model response (Response), which contribute to the KV cache. Top Row (Full KV Cache): All KV states are retained, achieving high accuracy (60.72%) but leading to Out-Of-Memory (OOM) errors. Middle Row (KV Cache Eviction): Prompt-related KV cache is compressed each turn. This reduces cache size by 50% but results in a severe accuracy drop to 17.33%. Bottom Row (FlowKV - Ours): Our proposed FlowKV method also reduces cache size by 50% by selectively compressing only the current turn’s prompt-related KV cache while preserving already compressed states from previous turns. This strategy effectively mitigates OOM issues while maintaining a high accuracy of 56.72%, significantly outperforming simple eviction.
2 FlowKV: Multi-Turn Context Isolation
2.1 Preliminaries: KV Cache Dynamics in Multi-Turn Conversations
The KV cache is a fundamental component in Transformer-based LLMs (Vaswani et al., 2017), designed to enhance the efficiency of autoregressive token generation. It stores the computed key and value vectors for all preceding tokens in a sequence, thereby obviating the need for their recomputation at each subsequent generation step.
In a multi-turn conversational setting, the context grows with each turn. Let $P_{sys}$ denote the initial system prompt. For turn $t$, let $Q_{t}$ be the user query and $R_{t}$ be the model’s response, $C_{t}$ be the KV cache pool at turn $t$. Under a standard Full KV Cache strategy, the entire conversational history is maintained. The KV cache accumulation can be described as follows:
The Initial State involves the system prompt.
$C_{0}=\text{KV}(P_{sys})$ (1)
For Turn 1, the model processes the first user query $Q_{1}$ using the context of $P_{sys}$. The KV cache for the prompt segment leading to the generation of $R_{1}$ is:
$C_{1}=\text{KV}(P_{sys}\oplus Q_{1})$ (2)
where $\oplus$ denotes sequence concatenation. After $R_{1}$ is generated, its KV pairs are added. The total KV cache state after Turn 1, $C_{1}$, becomes:
$C_{1}^{\prime}=C_{1}\oplus\text{KV}(R_{1})=\text{KV}(P_{sys}\oplus Q_{1}\oplus R_{1})$ (3)
This $C_{1}^{\prime}$ is carried forward.
For any subsequent Turn $t$ ($t>1$), let $H_{t-1}$ be the complete conversational history up to the end of turn $t-1$:
$H_{t-1}=P_{sys}\oplus Q_{1}\oplus R_{1}\oplus\dots\oplus Q_{t-1}\oplus R_{t-1}$ (4)
The KV cache state from the previous turn is $C_{t-1}=\text{KV}(H_{t-1})$. To generate $R_{t}$, the model processes $Q_{t}$ using the context $H_{t-1}$. The prompt KV cache for generating $R_{t}$ is:
$C_{t}=\text{KV}(H_{t-1}\oplus Q_{t})$ (5)
After $R_{t}$ is generated, the total KV cache state $C_{t}$ becomes:
$C_{t}^{\prime}=C_{t}\oplus\text{KV}(R_{t})=\text{KV}(H_{t-1}\oplus Q_{t}\oplus R_{t})$ (6)
This cumulative growth means $C_{t}$ includes all KV pairs from $P_{sys}$, all $Q_{i}$ and $R_{i}$ for $i\leq t$. While this Full KV Cache approach ensures maximum contextual information for the model, its linear growth with conversation length leads to substantial memory demands and computational overhead, presenting a critical challenge for deploying LLMs in extended dialogues.
In contrast to the Full KV Cache, many traditional KV cache eviction strategies aim to reduce memory by compressing the conversational history accumulated up to the current turn, before processing the current user query. A common approach, similar to mechanisms found in frameworks like NVIDIA’s kvpress (Jegou et al., 2024), while leaving the KV cache of the current turn’s query uncompressed for better performance. Let $\mathcal{F}(\cdot)$ denote a lossy KV cache compression.
At Turn 1: The KV cache of the system prompt, $P_{sys}$, is compressed:
$C_{1}=\mathcal{F}(\text{KV}(P_{sys}))\oplus\text{KV}(Q_{1})$ (7)
The response $R_{1}$ is then generated using the context formed by this compressed system prompt and the uncompressed KV cache of the first query. After $R_{1}$ is generated, the total KV cache state $C_{1}$ becomes:
$C_{1}^{\prime}=\mathcal{F}(\text{KV}(P_{sys}))\oplus\text{KV}(Q_{1}\oplus R_{1})$ (8)
At this point, the original $P_{sys}$ has been subjected to one compression operation.
At Turn 2: The $C_{1}^{\prime}$ will be compressed before the second turn response:
$C_{2}=\mathcal{F}(C_{1}^{\prime})\oplus\text{KV}(Q_{2})$ (9)
The response $R_{2}$ is generated using the context formed by this newly compressed history $C_{1}$ and the uncompressed KV cache of the second query.
As shown in Equation (9), the original system prompt information $P_{sys}$ (first compressed to $\mathcal{F}(P_{sys})$) is part of the input to a subsequent compression operation $\mathcal{F}(C_{1}^{\prime})$. Thus, $P_{sys}$ has effectively been compressed twice.
This pattern continues. After $T$ turns, the KV cache corresponding to the initial system prompt, $P_{sys}$, would have been subjected to $T$ nested compression operations on the path it contributes to the final compressed history. And $KV(Q_{i})$ and $KV(R_{i})$ for $i\leq t$ will be compressed $T-i$ times.
This repeated, nested compression of early conversational history is a primary cause of severe information degradation and context forgetting in such eviction strategies, leading to a rapid decline in coherence and task performance as the dialogue lengthens.
2.2 FlowKV Operational Mechanism
To address the challenge of information degradation from repeated compressions in naive KV cache eviction strategies (as discussed in Section 2.1), we introduce FlowKV. Designed as a multi-turn isolation mechanism, FlowKV’s core principle is to preserve the integrity of the already accumulated cache pool from previous turns ($C_{t-1}$), which contains previously compressed system prompts and uncompressed queries and responses. This approach mitigates catastrophic forgetting by avoiding re-compression of historical data. Futhermore, FlowKV is a general framework that can be integrated with any existing KV cache compression methods.
The operational mechanism of FlowKV is best understood by considering its behavior at each conversational turn $t$, and is visually contrasted with other methods in Figure 3 (Bottom Row).
Turn 1 (Identical to Traditional Eviction): For the first turn ($t=1$), FlowKV mirrors the behavior of the traditional eviction strategy described earlier.
$C_{1}=\mathcal{F}(\text{KV}(P_{sys}))\oplus\text{KV}(Q_{1})$ (10)
The response $R_{1}$ is then generated using the context formed by this compressed system prompt and the uncompressed KV cache of the first query. The total cache pool state after Turn 1, $C_{1}$, which is carried forward, incorporates the compressed system prompt, the uncompressed query, and the uncompressed response:
$C_{1}^{\prime}=\mathcal{F}(\text{KV}(P_{sys}))\oplus\text{KV}(Q_{1}\oplus R_{1})$ (11)
At Turn 2: When new inputs for the current turn query $Q_{2}$ is received, FlowKV focuses its compression effort on uncompressed part of $C_{1}$:
$C_{2}=\mathcal{F}{\color[rgb]{0.55859375,0.265625,0.6796875}(C_{1}^{\prime})}\oplus\text{KV}(Q_{2})$ (12)
Where ${\color[rgb]{0.55859375,0.265625,0.6796875}(C_{1}^{\prime})}$ is cache pool that apply FlowKV isolation mechanism, to be specific that the $\mathcal{F}(\text{KV}(P_{sys}))$ will preserved and only $KV(Q_{1}\oplus R_{1})$ will be compressed. Compare the equation (8) and (9), we can see that only the $P_{sys}$ is compressed twice, while $Q_{1}$ and $R_{1}$ are compressed once. With FlowKV, the $P_{sys}$, $KV(Q_{i})$ and $KV(R_{i})$ for $i\leq t$ will be have better preservation. This turn-by-turn isolation is the cornerstone of FlowKV. By not subjecting the accumulated history to repeated or joint compression, FlowKV minimizes the degradation of older contextual cues.
3 Experiments Design
3.1 Experimental Setups
This section will introduce the experimental setups, including the datasets, models, and evaluation environment.
Datasets
To evaluate the performance of LLMs under multi-turn conversation with different KV Cache compression strategies, we assess the Multi-IF He et al. (2024) and the PrefEval Zhao et al. (2025) datasets. Multi-IF focuses on evaluating the continuous instruction-following ability of LLMs across multiple turns of conversation, and directly identifies the influence of compression strategies on long-range context dependence and instruction execution accuracy. Besides, PrefEval provides diverse and challenging conversation prompts with hidden user preferences. It can be used to evaluate the impact of compression on the model’s ability to maintain conversation coherence and follow user preferences in complex interactions.
Evaluation Metrics
In terms of the Multi-IF dataset, we use Instruction Following Rate (IFR) to evaluate the model’s performance when facing multiple instructions:
$\text{IFR}=\frac{\text{SPA}+\text{SIA}+\text{LPA}+\text{LIA}}{4},$ (13)
where SPA refers to Strict Prompt-level Accuracy, SIA denotes Strict Instruction-level Accuracy, LPA means Loose Prompt-level Accuracy, and LIA signifies Loose Instruction-level Accuracy. Instruction-level calculates the percentage of followed instructions out of the total instructions within the same prompt, while Prompt-level assesses if the reply follows all instructions in the prompt, failing entirely if any single instruction is not followed. Strict adherence demands exact correspondence with the instruction, whereas Loose relaxes this by removing empty lines and insignificant symbols, among other things.
In terms of PrefEval, we use GPT-4o as a judge. For each model response, we provide GPT-4o with the response and the explicit user preference from the dataset and ask it if the response follows that preference. The User Preference Following Rate is then just the percentage of "yes" answers from GPT-4o.
Models
We conduct experiments on two LLMs, including LLaMA-3.1-8B-Instruct Grattafiori et al. (2024) and Qwen-2.5-7B-Instruct Yang et al. (2024a).
KV Cache Compression Methods
We selected the following KV cache compression methods for a comprehensive investigation of their effects: SnapKV Li et al. (2024b), StreamingLLM Xiao et al. (2023), ExpectedAttention Jegou et al. (2024), ChunkKV Liu et al. (2025b) and FlowKV. For below experiments, baseline means the method without FlowKV.
Evaluation Environments
We apply the kvpress Jegou et al. (2024) library to load models and KV Cache compression methods and evaluate their performance using a server with NVIDIA A40 GPUs. All experiments are conducted three times and the average results are reported.
3.2 Performance Comparison on Multi-IF
In this section, we will present and analyze the instruction following the performance of FlowKV and the baseline on the Multi-IF dataset.
[…]
Turn 1 Turn 2 $\Delta$ Turn 3 $\Delta$
LLaMA FKV - 73.41% 64.49% 56.62%
SKV Baseline 76.15% 37.08% 29.39%
FlowKV 76.15% 61.93%+24.85 54.95%+25.56
SLM Baseline 72.78% 33.94% 28.94%
FlowKV 72.78% 39.06%+5.12 41.58%+12.64
EA Baseline 76.05% 36.28% 30.48%
FlowKV 76.05% 64.89%+28.61 55.36%+24.88
CKV Baseline 70.47% 12.56% 16.49%
FlowKV 70.47% 52.83%+40.27 50.15%+33.66
Qwen FKV - 76.30% 60.72% 51.19%
SKV Baseline 76.49% 17.33% 21.96%
FlowKV 76.49% 56.72%+39.39 49.67%+27.71
SLM Baseline 76.47% 17.31% 21.08%
FlowKV 76.47% 36.82%+19.51 35.29%+14.21
EA Baseline 75.52% 17.62% 22.00%
FlowKV 75.52% 50.62%+33.00 39.25%+17.25
CKV Baseline 73.19% 18.47% 21.51%
FlowKV 73.19% 47.27%+28.80 43.55%+22.04
Table 1: Overall IFR Comparison on the Multi-IF Dataset. All results are reported with a KV compression ratio of 0.5. For FlowKV, delta is the absolute value. # FKV: FullKV, SKV: SnapKV, SLM: StreamingLLM, EA: ExpectedAttention, CKV: ChunkKV
As shown in Table 1, during the initial turn of the conversation, the core isolation mechanism of FlowKV is not yet engaged due to the absence of prior conversation history. However, starting from the second turn and continuing into the third turn, FlowKV outperforms the baseline as the contextual information accumulates and the compression mechanism begins to exert a significant effect. FlowKV achieves state-of-the-art IFR performance in the subsequent conversation turns for the current configuration. On average, there is a performance improvement of over 20%. With FlowKV, SnapKV and ExpectedAttention exhibit minimal performance degradation compared to FullKV, demonstrating that FlowKV’s multi-turn isolation strategy can maximally preserve information flow in multi-turn dialogues. For StreamingLLM, this method discards intermediate parts of the KV Cache, which inherently leads to the loss of intermediate information in multi-turn dialogues, resulting in significant performance degradation.
To gain a more comprehensive understanding of the robustness of FlowKV under different compression ratios, we further examined the performance of each method across a range of compression ratios from 0.1 to 0.9, where a higher compression ratio corresponds to a smaller KV cache size. As shown in Figure 4, we first compare the overall trends of the yellow line (SnapKV) and the green line (ExpectedAttention). It is evident that the performance of all methods decreases as the compression ratio increases, which is consistent with the expected information loss caused by higher compression. Next, by comparing FlowKV (triangle markers) with the baseline, we observe that FlowKV consistently outperforms the baseline in both Turn 2 and Turn 3. Notably, FlowKV achieves substantial performance improvements across all compression ratios. The gains are particularly pronounced at lower compression ratios (0.1–0.4), while at higher compression ratios (0.5–0.9), FlowKV still provides significant improvements, although the margin is relatively smaller compared to the low compression regime. These results indicate that FlowKV can robustly preserve the information flow across multiple dialogue turns, irrespective of the compression ratio.
3.3 Performance Comparison on PrefEval
Strategy SKV SLM EA CKV
LLaMA-3.1-8B-Instruct Full KV: 77.00% $\uparrow$
Baseline 10.60% 9.80% 10.90% 6.70%
FlowKV 58.70% 24.40% 75.40% 38.80%
Qwen-2.5-7B-Instruct Full KV: 55.90% $\uparrow$
Baseline 11.80% 11.60% 10.60% 10.30%
FlowKV 33.80% 16.80% 29.80% 26.40%
Table 2: Overall User Preference Following Rate Comparison on the PrefEval Dataset. All results are reported with a 0.5 KV compression ratio.
Figure 4: LLaMA-3.1-8B-Instruct on Multi-IF with different compression methods and strategies.
Figure 5: Case Study Comparing Multi-turn Strategies and KV Cache Compression for LLaMA-3.1-8B-Instruct on Multi-IF. Baseline represents SnapKV with 50% compression ratio, and FlowKV represents the FlowKV add on the Baseline. # [OMITTED] indicates portions of the text hidden for brevity.
The Table 2 shows the performance of different compression methods on PrefEval with a compression ratio of 0.5. Firstly, it can be seen that the baselines’ scores are very low comparing to the Full KV Cache, indicating that traditional KV Cache compression causes severe performance loss in multi-turn dialogue human preference benchmarks. When FlowKV is added, the performance improves significantly. For the LLaMA-3.1-8B-Instruct model with the ExpectedAttention method, the performance increased from 10.90% to 75.40%, an improvement of 64.5 percentage points. FlowKV also shows very significant performance improvements in Qwen-2.5-7B-Instruct. This indicates that the information flow preserved by FlowKV can effectively enhance the ability to maintain human preferences in multi-turn dialogues.
To further investigate the performance of FlowKV under varying compression strengths and its compatibility with different KV cache management methods, we plotted the user preference following rate as a function of the compression ratio. As shown in Figure 6, we present results for LLaMA-3.1-8B-Instruct with SnapKV and ExpectedAttention, both with and without FlowKV, on the PrefEval benchmark. First, by comparing the full KV cache and the baselines without FlowKV (lines without markers), we observe a substantial performance drop even at low compression ratios. This finding highlights that user preference following in multi-turn conversations is highly sensitive to the compression ratio. When FlowKV is incorporated, the performance improves markedly. At lower compression ratios (0.1–0.4), FlowKV is able to recover performance to a level close to that of the full KV cache. At higher compression ratios (0.5–0.9), FlowKV continues to provide significant improvements, although the gains are somewhat reduced compared to the low compression ratio. These results demonstrate that FlowKV can robustly preserve information flow across multiple dialogue turns, regardless of the compression ratio. Furthermore, our findings reveal that user preference following in multi-turn dialogue scenarios remains a highly challenging task.
Figure 6: LLaMA-3.1-8B-Instruct on PrefEval with different compression methods and strategies.
3.4 Case Study
To better illustrate the impact of model generation quality across different multi-turn strategies and KV Cache compression methods, we conduct a case study to demonstrate the performance differences of the LLaMA-3.1-8B-Instruct model when processing multi-turn cumulative instructions on the Multi-IF dataset. Figure 5 shows the Full KV Cache setting successfully follows all instructions of each turn, accurately generating content that satisfies format control, style restrictions, stipulated texts, and other complex tasks. However, the Baseline setting leads to a significant performance crash: not only does content duplication occur in the second turn, violating the newly added "p.p.s" instructions, but in the third turn, the "p.p.s" and double quotation instructions fail to be fully executed, exposing severe context forgetting and the loss of processing capabilities for the conversation flow. When using FlowKV multi-turn conversation, the model shows relative adaptability and performance trade-offs. Although the requirement of all-lowercase format accumulated in the first turn is not fully maintained from the second turn, the content generation still maintains relevance and creativity to the topic and successfully executes the "p.p.s" marks and double quotation marks wrapping instructions newly added in the second and third turns, respectively. This indicates that FlowKV has effectively preserved historical information flow also keep the low KV Cache compression ratio. More case study can be found in Appendix C.
4 Related Works
LLMs recently made significant progress in text understanding and content generation. However, LLMs face challenges when handling multi-turn interactions. Hence, two research fields attract much attention for improving the efficiency and feasibility in real practice: KV Cache optimization on LLMs and multi-turn conversation in LLMs.
4.1 Key–Value Cache Optimization Strategies on LLMs
The attention mechanism needs to compute the relationship between the current token and previous tokens when autoregressive generation. LLMs introduce KV Cache to avoid re-computing the Key and Value vectors of previous tokens. However, the size of KV Cache has a linear relationship with the sequence length Bai et al. (2023), and it may increase sharply when handling long context or multi-turn conversation. To alleviate the challenges brought by KV Cache, researchers proposed multiple compression and optimization strategies, which can be roughly divided into quantization, sparsification (pruning and eviction), low-rank approximation, and cross-layer sharing. The token sparsification here identifies and selectively retains the KV Cache corresponding to the most important tokens, discarding the relatively unimportant parts. For instance, H2O Zhang et al. (2023) considers the token with a higher cumulative attention score to be more important, while FastGen Ge et al. (2023) customizes different strategies for heads by analyzing the behavioral patterns of attention heads. StreamingLLM Xiao et al. (2023) retains only some of the initial tokens and a fixed-size window of the most recent tokens based on its discovery that these tokens are crucial for maintaining the model’s performance. In addition, SnapKV Li et al. (2024b) compresses KV caches by selecting key attention head features from a prompt, enabling efficient long-sequence processing with minimal performance loss. ChunkKV Liu et al. (2025b) compresses KV caches by grouping tokens into semantic chunks, preserving contextual meaning while reducing memory usage. Besides, low-rank approximation Saxena et al. (2024) and cross-layer sharing Sun et al. (2024); Brandon et al. (2024); Liu et al. (2024) and others are also important optimization directions.
4.2 Multi-turn Conversation in LLMs
Multi-turn conversation represents a crucial application scenario for LLMs, simulating human conversation to understand context and generate coherent, relevant responses over successive interaction turns Zhang et al. (2019); Shuster et al. (2022); Thoppilan et al. (2022); Wu et al. (2020). However, as the number of turns increases, LLMs encounter challenges. Current attention is largely focused on Long Context Handling Yi et al. (2024), which involves treating preceding turns as contextual information and managing the growing conversation history, especially given the constraints of models’ limited context windows. Furthermore, Coherence and Consistency Li et al. (2025) pose another challenge, concerning the need to ensure that responses generated in later turns remain consistent with earlier parts of the conversation and avoid contradictions.
5 Conclusion
In this paper, we addressed the critical challenge of KV Cache management in LLMs for multi-turn conversational scenarios, where the trade-off between computational efficiency and contextual coherence is paramount. We identified that conventional KV Cache eviction strategies often lead to severe information degradation due to the repeated compression of historical context. To overcome this limitation, we introduced FlowKV, a novel multi-turn isolation mechanism. FlowKV’s core principle is to safeguard the integrity of the already compressed KV cache from prior turns, strategically applying compression only to the newly generated segments of the most recent turn. This training-free approach, compatible with existing KV compression techniques, effectively mitigates catastrophic forgetting and preserves vital long-range dependencies.
Limitations
FlowKV’s performance, while significantly enhancing multi-turn coherence, is intrinsically tied to the efficacy of the base KV cache compression algorithm applied to the current turn’s data. Furthermore, additional experiments on a broader range of tasks, such as multi-turn coding, math, and reasoning dialogues, would be beneficial to more comprehensively evaluate FlowKV’s effectiveness and generalization.
Acknowledgments
This work was supported by Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
Appendix A Additional Results
A.1 Efficiency Results
To evaluate the efficiency of FlowKV, we conduct experiments in a fixed prompt and output length setting, as summarized in Table 3. We compare three configurations: FullKV (baseline without compression), ChunkKV, and ChunkKV integrated with FlowKV. As shown in the table, FlowKV does not compromise the efficiency of existing KV cache compression techniques. It maintains a high compression ratio (0.9) comparable to ChunkKV alone, while introducing negligible overhead in prefilling time and total generation time. Moreover, FlowKV achieves similar performance in terms of TTFT (Time to First Token) and TPOT (Time Per Output Token), demonstrating that it is an efficient strategy that incurs no significant additional computational cost. These results confirm that FlowKV can be seamlessly integrated with KV cache compression methods to further enhance system efficiency without sacrificing speed or resource usage.
Table 3: Efficiency Results for FlowKV
Configuration Prompt Output Compression Prefilling Cache TTFT TPOT Total Gen.
Length Length Ratio Time(s) $\downarrow$ Size(GB) $\downarrow$ (s) $\downarrow$ (ms) $\downarrow$ Time(s) $\downarrow$
FullKV 8192 4096 - 1.5621 1.0000 1.6013 45.9421 184.2934
ChunkKV 8192 4096 0.9 1.3653 0.1000 1.3914 39.8702 164.6600
ChunkKV+FlowKV 8192 4096 0.9 1.3632 0.1000 1.4002 39.8812 165.2100
Figure 7 and 8 provides a detailed turn-by-turn breakdown of the Instruction Following Rate (IFR) for LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct on the Multi-IF dataset. It compares the performance of four KV cache compression methods (SnapKV, StreamingLLM, ExpectedAttention, and ChunkKV) when integrated with either the FlowKV strategy or a baseline approach. The subplots distinctly illustrate performance at Turn 2 and Turn 3 across a spectrum of compression ratios, with the Full KV Cache performance serving as a reference (red dashed line). This visualization highlights how different compression methods interact with FlowKV and the baseline strategy as conversational depth increases and compression becomes more aggressive.
Figure 9 illustrates the User Preference Following Rate for LLaMA-3.1-8B-Instruct on the PrefEval dataset. The figure contrasts the FlowKV strategy against the baseline strategy across various compression ratios, for different underlying KV cache compression methods (SnapKV, StreamingLLM, ExpectedAttention, and ChunkKV). The performance of a Full KV Cache (uncompressed) is shown as a dashed red line reference. This visualization allows for an assessment of how well user preferences are maintained under increasing compression levels when FlowKV’s isolation mechanism is employed compared to a standard baseline approach.
Figure 7: LLaMA3.1-8B-Instruct Multi-IF Instruction Following Performance.
Figure 8: Qwen-2.5-7B-Instruct Multi-IF Instruction Following Performance.
Figure 9: Performance Comparison of Multi-turn Conversation Strategies on LLaMA-3.1-8B-Instruct Using PrefEval Dataset.
The following table 4, table 6, and table 5 provide comprehensive numerical results complementing the figures. They detail the performance of both LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models on the Multi-IF and PrefEval datasets. For the Multi-IF dataset, Instruction Following Rates (IFR) are presented turn-by-turn (Turn 1, Turn 2, and Turn 3). For the PrefEval dataset, the User Preference Following Rate is reported. All results are shown for different KV cache compression methods (SnapKV, StreamingLLM, ExpectedAttention, ChunkKV), comparing the FlowKV strategy against the baseline approach, across a fine-grained spectrum of compression ratios ranging from 0.1 to 0.9. The performance with a Full KV Cache is also provided as a key reference point for each model and turn where applicable.
KV Method Strategy 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Turn 1
LLaMA-3.1-8B-Instruct Full KV: 73.41%
SKV Baseline 74.85% 76.53% 74.81% 73.77% 76.15% 75.97% 74.06% 75.57% 76.98%
FlowKV 74.85% 76.53% 74.81% 73.77% 76.15% 75.97% 74.06% 75.57% 76.98%
SLM Baseline 73.17% 75.59% 75.31% 75.07% 72.78% 72.86% 73.29% 73.41% 76.39%
FlowKV 73.17% 75.59% 75.31% 75.07% 72.78% 72.86% 73.29% 73.41% 76.39%
EA Baseline 74.85% 75.00% 75.04% 74.31% 76.05% 75.57% 75.23% 75.62% 76.35%
FlowKV 74.85% 75.00% 75.04% 74.31% 76.05% 75.57% 75.23% 75.62% 76.35%
CKV Baseline 74.80% 74.80% 74.80% 70.47% 70.47% 70.47% 70.49% 70.49% 70.50%
FlowKV 74.80% 74.80% 74.80% 70.47% 70.47% 70.47% 70.49% 70.49% 70.50%
Turn 2
LLaMA-3.1-8B-Instruct Full KV: 64.49%
SKV Baseline 35.99% 34.89% 35.62% 38.09% 37.08% 35.30% 36.65% 34.39% 19.09%
FlowKV 65.26% 65.62% 63.23% 62.54% 61.93% 58.72% 54.84% 49.13% 29.28%
SLM Baseline 35.21% 34.96% 36.44% 36.99% 33.94% 33.29% 33.46% 32.32% 31.09%
FlowKV 64.44% 56.74% 48.42% 41.96% 39.06% 35.20% 29.77% 27.04% 25.35%
EA Baseline 36.33% 35.79% 36.37% 36.20% 36.28% 35.49% 35.04% 34.14% 31.08%
FlowKV 66.17% 65.51% 64.88% 64.56% 64.89% 64.05% 63.59% 60.42% 53.12%
CKV Baseline 35.72% 35.68% 35.74% 12.54% 12.56% 12.58% 12.56% 12.54% 12.55%
FlowKV 64.82% 63.54% 61.49% 57.82% 52.83% 45.49% 38.06% 30.23% 24.01%
Turn 3
LLaMA-3.1-8B-Instruct Full KV: 56.62%
SKV Baseline 30.30% 29.81% 30.84% 30.47% 29.39% 29.89% 29.76% 28.00% 17.40%
FlowKV 57.06% 56.78% 56.41% 55.59% 54.95% 52.21% 50.20% 43.84% 30.92%
SLM Baseline 29.82% 30.43% 31.16% 31.52% 28.94% 27.50% 28.11% 28.12% 28.26%
FlowKV 55.66% 54.40% 50.78% 46.65% 41.58% 37.19% 31.60% 26.28% 23.54%
EA Baseline 30.33% 29.93% 30.93% 30.02% 30.48% 29.50% 30.16% 29.68% 28.24%
FlowKV 56.48% 57.28% 57.00% 56.99% 55.36% 56.33% 54.75% 52.21% 47.24%
CKV Baseline 30.27% 30.28% 35.77% 16.47% 16.49% 16.47% 16.49% 16.50% 16.47%
FlowKV 56.71% 55.80% 55.58% 53.52% 50.15% 44.77% 38.11% 29.55% 23.05%
Table 4: LLaMA-3.1-8B-Instruct Performance on Multi-IF across Different KV Cache Methods and Ratios
KV Method Strategy 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Turn 1
Qwen-2.5-7B-Instruct Full KV: 76.30%
SKV Baseline 77.52% 76.96% 76.36% 75.77% 76.49% 72.62% 50.78% 49.23% 71.82%
FlowKV 77.52% 76.96% 76.36% 75.77% 76.49% 72.62% 50.78% 49.23% 71.82%
SLM Baseline 77.20% 77.12% 76.55% 76.70% 76.47% 77.15% 76.36% 73.59% 72.75%
FlowKV 77.20% 77.12% 76.55% 76.70% 76.47% 77.15% 76.36% 73.59% 72.75%
EA Baseline 76.39% 77.05% 77.59% 77.24% 75.52% 75.05% 75.02% 73.64% 72.85%
FlowKV 76.39% 77.05% 77.59% 77.24% 75.52% 75.05% 75.02% 73.64% 72.85%
CKV Baseline 73.14% 73.28% 73.19% 73.19% 73.19% 73.28% 73.28% 73.23% 73.23%
FlowKV 73.14% 73.28% 73.19% 73.19% 73.19% 73.28% 73.28% 73.23% 73.23%
Turn 2
Qwen-2.5-7B-Instruct Full KV: 60.72%
SKV Baseline 16.70% 15.93% 17.59% 17.35% 17.33% 18.29% 17.24% 17.20% 20.23%
FlowKV 61.24% 60.17% 58.95% 58.73% 56.72% 53.59% 45.50% 44.48% 36.81%
SLM Baseline 16.77% 16.39% 17.23% 17.13% 17.31% 17.66% 18.66% 17.77% 19.35%
FlowKV 58.05% 51.67% 45.64% 42.63% 36.82% 33.83% 30.31% 28.46% 26.48%
EA Baseline 16.42% 16.57% 17.39% 17.14% 17.62% 17.99% 17.72% 17.80% 19.34%
FlowKV 59.54% 60.46% 57.48% 54.49% 50.62% 47.08% 41.26% 36.65% 31.26%
CKV Baseline 18.47% 18.46% 18.47% 18.46% 18.47% 18.47% 18.48% 18.48% 18.47%
FlowKV 59.52% 57.18% 53.82% 50.99% 47.27% 42.68% 37.25% 29.78% 25.03%
Turn 3
Qwen-2.5-7B-Instruct Full KV: 51.19%
SKV Baseline 21.27% 20.77% 21.83% 21.77% 21.96% 21.84% 18.83% 17.98% 22.35%
FlowKV 51.33% 50.97% 49.41% 50.07% 49.67% 47.11% 41.61% 39.74% 32.41%
SLM Baseline 21.03% 20.57% 20.86% 20.80% 21.08% 21.86% 22.06% 22.02% 22.90%
FlowKV 47.86% 44.40% 42.90% 39.55% 35.29% 32.20% 30.01% 26.53% 22.06%
EA Baseline 21.77% 21.16% 21.94% 20.93% 22.00% 21.34% 21.78% 22.06% 22.90%
FlowKV 50.84% 49.74% 47.11% 43.69% 39.25% 35.22% 30.02% 24.89% 21.75%
CKV Baseline 21.50% 21.48% 21.50% 21.49% 21.51% 21.48% 21.51% 21.52% 21.50%
FlowKV 49.49% 49.19% 47.08% 46.76% 43.55% 40.01% 35.96% 29.32% 22.68%
Table 5: Qwen-2.5-7B-Instruct Performance on Multi-IF across Different KV Cache Methods and Ratios
KV Method Strategy 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
LLaMA-3.1-8B-Instruct Full KV: 77.00%
SKV Baseline 18.00% 15.70% 12.30% 12.60% 10.60% 8.80% 8.00% 9.10% 6.40%
FlowKV 77.30% 75.40% 70.70% 64.90% 58.70% 49.40% 39.50% 30.90% 20.40%
SLM Baseline 17.90% 16.20% 12.90% 11.40% 9.80% 7.80% 7.60% 6.30% 6.70%
FlowKV 76.20% 67.00% 39.30% 29.50% 24.40% 21.40% 19.20% 18.10% 12.70%
EA Baseline 18.50% 16.40% 13.60% 12.50% 10.90% 10.20% 9.30% 10.00% 6.50%
FlowKV 77.90% 76.90% 78.00% 76.00% 75.40% 71.00% 69.50% 59.60% 49.40%
CKV Baseline 12.60% 13.10% 8.40% 8.40% 6.70% 6.40% 6.70% 6.40% 6.20%
FlowKV 76.30% 71.80% 62.10% 49.90% 38.80% 29.10% 22.90% 18.50% 13.70%
Table 6: LLaMA-3.1-8B-Instruct Performance on PrefEval across Different KV Cache Methods and Ratios
Appendix B Additional Related Work
B.1 Multi-Turn Conversation
Long-term Dependency Sun et al. (2025) is also a significant focus, requiring models to accurately grasp entities, relationships, and user intentions established early in the conversation and generate responses based on this understanding. To address these challenges, integrating Retrieval-Augmented Generation (RAG) Lewis et al. (2020) with LLMs can improve the accuracy and relevance of responses. Additionally, Yi et al. (2024) and Madotto et al. (2019) mention that explicitly modeling the structural information of the conversation can effectively convey inter-turn dependencies and speaker information to the model for reference. Benchmark datasets for multi-turn conversation often evaluate the ability of LLMs to follow instructions or retain information over multiple accumulated turns. For instance, He et al. (2024) introduces the Multi-IF dataset to assess the degree to which different LLMs continuously adhere to user instructions across turns in multi-turn scenarios. Zhao et al. (2025) constructs the PrefEval dataset, containing user preference and query pairs presented in a multi-turn format, designed to evaluate the capability of large LLMs to infer, remember, and follow user preferences within long-context, multi-turn conversations. SCBench (Li et al., 2024a) is a comprehensice benchmark for evaluating long-context, and multi-turn conversation is main subtask.
Appendix C Case Study
Figure 10: Case Study Comparing Multi-Turn Strategies and KV Cache Compression for LLaMA-3.1-8B-Instruct on PrefEval. Full KV Cache represents no compression, FlowKV represents the FlowKV multi-turn strategy with 50% SnapKV, and Baseline represents the Baseline multi-turn strategy with 50% SnapKV. User preference is hidden in the historical conversation. The model should respond to the final query according to inferred user preference.
In addition, we provide another case study for the PrefEval dataset as Figure 10. Different from Multi-IF, PrefEval requires the model to infer the user preference according to the provided historical conversation, which brings a greater challenge. The case study is conducted on Full KV Cache, 50% SnapKV FlowKV multi-turn conversation, and 50% SnapKV Baseline multi-turn conversation. The user’s preference of preferring peer-to-peer interaction is hidden in the previous conversations; for example, the user mentions she prefers learning in group settings. The final query is to make the model recommend resources for data analysis learning. By successfully inferring the user preference, Full KV strategy and FlowKV provide some group-based learning resources. Moreover, the resources provided by Full KV are more authentic. This indicates that FlowKV has effectively preserved historical information but still suffers from some performance losses. However, the Baseline, which is 50% SnapKV Baseline multi-turn conversation here, completely failed to infer the hidden user preferences. Instead, it tended to answer the user’s first question from the conversation history. This indicates that in long-context situations, the impact of KV compression is devastating, potentially leading to the generation of unknown responses. Because the conversation history of PrefEval is significantly longer than that of Multi-IF, the performance decline of Baseline will be more pronounced compared to Multi-IF.
Appendix D Evaluation Benchmark
D.1 Dataset Details
Detailed statistics for each benchmark dataset are provided in Table 7.
DATASET TASK TYPE # TEST METRIC
Multi-IF He et al. (2024) Multi-Turn Instruction Following 909 Instruction Following Rate (%)
PrefEval Zhao et al. (2025) Multi-Turn Preference Inference 1000 Preference Following Rate (%)
Table 7: The statistics of the datasets used in this paper.# Test denotes the number of test data.
Appendix E Licenses
For the evaluation dataset, Multi-IF (He et al., 2024) is released under Apache-2.0 license. PrefEval (Zhao et al., 2025) is released under the Attribution-NonCommercial 4.0 International license.